### v1

@@ 1 Introduction
The proximal operator is a key mathematical tool in solving regularised problems of the form given in (2). For a function $R(x)$, its proximal operator is defined as:
@@
Building on this foundation, PnP has evolved significantly since its introduction by [7]. Early approaches demonstrated its effectiveness across diverse imaging tasks [8], while subsequent studies [9, 10, 11] refined the framework by enhancing stability, convergence, and applicability. Additionally, works such as [12, 13, 14, 15, 10] extended its reach to more complex imaging scenarios. A notable milestone was the development of TFPnP (Tuning-Free Plug-and-Play) [16], which eliminated the need for parameter tuning, simplifying practical deployment.
The core of the Plug-and-Play (PnP) framework lies in the denoiser, $H_{\sigma}$, which serves as an implicit regulariser for the reconstruction process. This denoiser is typically trained using a large amount of data, allowing it to learn rich priors that generalise well to diverse instances of the underlying task. PnP frameworks have been developed utilising a variety of denoisers. Classical methods like BM3D [17] and TV [18] remain popular due to their robustness and minimal reliance on training data. In the meantime, the rise of deep learning-based denoisers [8, 14, 19] has significantly advanced the field by capturing complex data patterns, such as DnCNN [20] or U-Net [21] achieve their performance by leveraging extensive datasets to capture the statistical properties of natural images or specific domains. However, this reliance on large-scale training datasets poses a significant challenge in domains where data availability is limited.
@@
We present the first framework to incorporate Kolmogorov-Arnold Networks (KANs) as denoisers in the PnP paradigm, focusing on solving inverse problems with single-instance priors:
@@ 2 Proposed Method
$\min_{x,z}f(x)+\lambda R(H_{\sigma}(z)),$ (5)
where $f(x)$ is the data fidelity term, $\lambda$ is a regularisation parameter, and $R(x)$ is the regularisation functional implicitly defined through a network denoiser $H_{\sigma}(x)$. This framework generalises a wide range of inverse problems.
@@ 2.1 KAN as the Ideal Single-Instance Prior
In our framework, we focus on a single-shot learning setting, where we only have a single noisy observation of the underlying clean signal. This presents a challenge, as traditional denoisers may struggle in such settings, where the noise characteristics are not fixed across samples. To address this, we propose a novel prior drawing from Kolmogorov-Arnold Networks (KANs) as the denoiser. KANs are a powerful class of neural networks based on the Kolmogorov-Arnold representation theorem, which ensures that any continuous function can be approximated by a finite composition of smooth, univariate functions. This ability to smoothly approximate complex mappings makes KANs particularly well-suited for denoising tasks, as they naturally preserve the underlying structure of the signal while removing noise. Kolmogorov-Arnold Networks (KANs) are a class of neural networks based on the Kolmogorov-Arnold representation theorem, which guarantees that any continuous function on a multidimensional domain can be approximated by a finite composition of continuous functions. KAN is defined as follows. Let $x\in\mathbb{R}^{d}$ be the input vector, and let $y\in\mathbb{R}$ be the output. A KAN approximates a continuous function $f(x)$ by composing several layers of univariate basis functions. Specifically, a KAN can be expressed as:
@@ 2.2 Key Assumptions for KAN-PnP
This implies that $R(x)$ is bounded below due to the quadratic penalty term, which prevents $R(x)$ from decreasing unboundedly. However, in the single-shot setting, the behavior of $R(x)$ depends on how $H_{\sigma}$ is constructed and adapted. To ensure boundedness:
@@
Differentiability of the Implicit Regulariser. In gradient-based formulations, the implicit regulariser $R(x)$ derived from $H_{\sigma}$ must have a Lipschitz continuous gradient. This condition can be satisfied if $H_{\sigma}(x)$ is differentiable and smooth, which follows from the construction of the KAN denoiser using B-splines. In single-shot learning, we ensure the smoothness of $H_{\sigma}$, avoiding overfitting artifacts or discontinuities introduced during adaptation to the single image.
@@ 2.3 Convergence of KAN-PnP with Single-Instance Prior
$\min_{x,z}f(x)+\lambda\cdot R(H_{\sigma}(z))\quad\text{subject to}\quad x=z,$
@@
    Lipschitz continuity of the KAN denoiser $H_{\sigma}$, which guarantees that the denoising step is non-expansive and stable.
@@
    Boundedness of the regularisation functional $R(x)$, ensuring that the regulariser does not diverge as the iterates grow large.
  4. 4.
@@
    Differentiability of the regulariser $R(x)$, if required for gradient-based updates, which is satisfied under the smoothness of $H_{\sigma}$.
Theorem 2.3 (Convergence of KAN-PnP).
Consider the optimisation problem
$\min_{x,z}f(x)+\lambda R(H_{\sigma}(z)),$
where $f(x)$ is the data fidelity term, $\lambda$ is a regularisation parameter, and $R(x)$ is the regularisation functional defined implicitly through the KAN denoiser $H_{\sigma}(x)$. The KAN-PnP with ADMM proceeds by iteratively solving the following steps of that (17). Given the assumptions that the KAN denoiser $H_{\sigma}$ is Lipschitz continuous, the data fidelity term $f(x)$ is convex, and the regularisation functional $R(x)$ is bounded below, the iterates $x^{k}$, $z^{k}$, and $u^{k}$ generated by the PnP-ADMM algorithm converge to a stationary point of the original optimisation problem.
Proof 2.4.
We begin by analysing the primal and dual residuals, which are commonly used in ADMM convergence analysis. The primal residual measures the inconsistency between $x^{k+1}$ and $z^{k+1}$:
$r^{k+1}=x^{k+1}-z^{k+1}.$
The dual residual measures the consistency of the dual updates:
$s^{k+1}=\mu(z^{k+1}-z^{k}).$
For convergence, we require that both the primal residual $r^{k+1}$ and the dual residual $s^{k+1}$ converge to zero as $k\to\infty$. Given that $H_{\sigma}$ is Lipschitz continuous and the data fidelity term $f(x)$ is convex, it follows that both residuals will converge to zero, provided the iterates remain bounded.
The augmented Lagrangian for the problem is given by:
$\mathcal{L}_{\mu}(x,z,u)=f(z)+\lambda R(H_{\sigma}(z))+\frac{\mu}{2}\|x-z+u\|^{2}.$
As established in standard ADMM convergence theory, if the iterates $x^{k}$, $z^{k}$, and $u^{k}$ satisfy the update rules, the augmented Lagrangian decreases monotonically over iterations:
$\mathcal{L}_{\mu}(x^{k+1},z^{k+1},u^{k+1})\leq\mathcal{L}_{\mu}(x^{k},z^{k},u^{k})\quad\text{for all}\quad k.$
This monotonic decrease of the augmented Lagrangian ensures that the objective function $f(x)+\lambda R(H_{\sigma}(z))$ converges to a minimum, and that the iterates $x^{k}$, $z^{k}$, and $u^{k}$ approach a stationary point of the given optimisation problem.
As the ADMM iterations progress, the iterates $x^{k}$, $z^{k}$, and $u^{k}$ converge to a stationary point that satisfies the first-order optimality conditions of the optimisation problem. Specifically, at the limit point, the following conditions hold:
$0\in\partial f(z^{*})+\mu(z^{*}-x^{*}-u^{*}),$
$0\in\partial R(H_{\sigma}(x^{*}))+\mu(x^{*}-z^{*}+u^{*}),$
where $\partial f(z)$ and $\partial R(H_{\sigma}(x))$ denote the subdifferentials of $f(z)$ and $R(H_{\sigma}(x))$, respectively. These conditions guarantee that the limit point is a stationary solution to the optimisation problem.
Thus, under the assumptions of Lipschitz continuity of $H_{\sigma}$, convexity of $f(x)$, and boundedness of $R(x)$, the PnP-ADMM algorithm converges to a stationary point of the optimisation problem, completing the proof.
@@ 3 Experimental Results & Discussion
For the experimental comparison, we followed the dual resizing strategies to preprocess the image. The images are sourced with Creative Commons Licenses which are resized to $512\times 384$, and the selected data in [26, 27] were tested. The experiment were conducted on NVIDIA A100 GPU with 80GB RAM. The Plug-and-Play phase ensued with the $\bigtriangledown$-Prox toolbox [28], adopting the default settings for all tasks.
@@
Evaluation Protocol. To ensure a robust comparison, we evaluated the Noise2Self pre-training approach [29] on three networks, each undergoing $100$ iterations with a learning rate of $0.01$. The FFDNet [30] model is designed with $8$ hidden layers and $64$ feature maps per layer, while the UNet [21] architecture featured a $4$-level downsampling scheme and began with $32$ feature maps in its initial convolutional layer. For consistency, the optimisation setup mirrored the configuration used in our proposed approach. We compare these results against both the Noise2Self approach and implicit neural MLP, as well as our proposed implicit KAN method. The evaluation of the methods were measured by Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM), where PSNR and SSIM scores signal are higher when reconstruction quality improved.
@@ 3.1 KAN-PnP in High-Stakes Image Reconstruction
Our proposed KAN-PnP framework has been extensively evaluated in two challenging inverse problem settings: super-resolution and joint optimisation for demosaicing and deconvolution. Super-resolution tests the model’s ability to enhance image detail across varying upscaling factors (2$\times$, 4$\times$ and 8$\times$), while the joint optimisation task emphasises its robustness and adaptability by tackling the intertwined challenges of demosaicing and deconvolution. The latter setting is particularly demanding due to the need to simultaneously address two operations—removing the effects of mosaic patterns while restoring the high-frequency details lost during image acquisition—requiring a balance of accuracy and computational stability.
@@ 4 Conclusion
In this work, we introduced KAN-PnP, an optimisation framework that integrated Kolmogorov-Arnold Networks (KANs) as denoisers within the Plug-and-Play (PnP) paradigm. KAN-PnP was specifically designed to tackle inverse problems with single-instance priors, where only a single noisy observation was available, thus eliminating the need for large datasets typically required by traditional denoising methods. We demonstrated that KANs, based on the Kolmogorov-Arnold representation theorem, effectively served as priors in these settings, offering a robust solution for denoising. We provided theoretical guarantees for KAN-PnP, proving its convergence under key assumptions. Extensive experimental results showed that KAN-PnP outperformed existing denoising methods, providing superior performance in single-shot learning scenarios with minimal data. The method exhibited reliable convergence, achieving high accuracy with fewer iterations. Comparisons with state-of-the-art approaches further highlighted KAN-PnP’s advantages in both convergence speed and solution quality, demonstrating its potential for solving real-world inverse problems with limited data.

### v2

@@ 1 Introduction
The proximal operator is a key mathematical tool [6] in solving regularised problems of the form given in (2). For a function $R(x)$, its proximal operator is defined as:
@@
Building on this foundation, PnP has evolved significantly since its introduction by [8]. Early approaches demonstrated its effectiveness across diverse imaging tasks [9], while subsequent studies [10, 11, 12, 13] refined the framework by enhancing stability, convergence, and applicability. Additionally, works such as [14, 15, 16, 17, 11] extended its reach to more complex imaging scenarios. A notable milestone was the development of TFPnP (Tuning-Free Plug-and-Play) [18], which eliminated the need for parameter tuning, simplifying practical deployment.
The core of the Plug-and-Play (PnP) framework lies in the denoiser, $H_{\sigma}$, which serves as an implicit regulariser for the reconstruction process. This denoiser is typically trained using a large amount of data, allowing it to learn rich priors that generalise well to diverse instances of the underlying task. PnP frameworks have been developed utilising a variety of denoisers. Classical methods like BM3D [19] and TV [20] remain popular due to their robustness and minimal reliance on training data. In the meantime, the rise of deep learning-based denoisers [9, 16, 21] has significantly advanced the field by capturing complex data patterns, such as DnCNN [22] or U-Net [23] achieve their performance by leveraging extensive datasets to capture the statistical properties of natural images or specific domains. Recent works on Weakly Convex Ridge Regularisers (WCRR-NNs) [24] proposes learning explicit regularisation functionals $R(x)$ such that the inverse problem remains (weakly) convex. These methods operate in a variational setting and typically require access to a dataset for training. However, this reliance on large-scale training datasets poses a significant challenge in domains where data availability is limited.
@@
We present the first framework to incorporate Kolmogorov-Arnold Networks (KANs) [28] as denoisers in the PnP paradigm, focusing on solving inverse problems with single-instance priors:
@@ 2 Proposed Method
$\min_{x,z}f(x)+\lambda R(H_{\sigma}(z)),\quad\text{subject to}\quad x=z$ (5)
where $f(x)$ is the data fidelity term, $\lambda$ is a regularisation parameter, and $R(x)$ is the regularisation functional implicitly defined through a network denoiser $H_{\sigma}(x)$. This framework generalises a wide range of inverse problems. We note that (5) should be understood as a conceptual variational formulation underpinning the Plug-and-Play scheme. While $R(H_{\sigma}(z))$ is not used explicitly in the algorithm, it allows us to interpret the denoising step as a surrogate proximal operator of an implicit regulariser $R$.
@@ 2.1 KAN as the Ideal Single-Instance Prior
In our framework, we focus on a single-shot learning setting, where we only have a single noisy observation of the underlying clean signal. This presents a challenge, as traditional denoisers may struggle in such settings, where the noise characteristics are not fixed across samples. To address this, we propose a novel prior drawing from Kolmogorov-Arnold Networks (KANs) as the denoiser. KANs are a powerful class of neural networks based on the Kolmogorov-Arnold representation theorem, which ensures that any continuous function can be approximated by a finite composition of smooth, univariate functions. This ability to smoothly approximate complex mappings makes KANs particularly well-suited for denoising tasks, as they naturally preserve the underlying structure of the signal while removing noise. KANs introduce a form of structured inductive bias through kernel-based interpolation over learnable anchors, which induces local smoothness and adaptivity — properties that are well-aligned with the statistical structure of images. Unlike MLP-based INRs, which are globally connected and prone to spectral bias (favoring low-frequency content), KAN’s piece- wise smooth interpolation better captures high-frequency details and edges, which are essential for image restoration tasks (imaging inverse problems). In the single- shot setting, where no external supervision is available, this inductive bias acts as a strong prior that regularises the learning dynamics and helps avoid overfitting to noise. Kolmogorov-Arnold Networks (KANs) are a class of neural networks based on the Kolmogorov-Arnold representation theorem [30, 31, 32], which guarantees that any continuous function on a multidimensional domain can be approximated by a finite composition of continuous functions. KAN is defined as follows. Let $x\in\mathbb{R}^{d}$ be the input vector, and let $y\in\mathbb{R}$ be the output. A KAN approximates a continuous function $f(x)$ by composing several layers of univariate basis functions. Specifically, a KAN can be expressed as:
@@
On the Theorem 1 and its Implications. While Lipschitz continuity of neural networks is well-known, we emphasise that our result is tailored to the Kolmogorov-Arnold Network (KAN) architecture, which differs significantly from standard ReLU-based networks. The use of smooth univariate B-spline basis functions requires a more explicit analysis to ensure global Lipschitz continuity, especially in the context of single-instance priors, where no data-driven regularisation is available. Although the derived Lipschitz bound in Theorem 1 may be conservative, it offers a principled way to control the smoothness and stability of the KAN denoiser via architectural parameters (e.g., number of grid points $G$, spline smoothness $k$, and layer width). This enables practitioners to design networks with desirable properties, including potential non-expansiveness ($L_{H}\leq 1$), which is beneficial for convergence in Plug-and-Play schemes.
@@ 2.2 Key Assumptions for KAN-PnP
This implies that $R(x)$ is bounded below due to the quadratic penalty term, which prevents $R(x)$ from decreasing. However, in the single-shot setting, the behavior of $R(x)$ depends on how $H_{\sigma}$ is constructed and adapted. To ensure boundedness:
@@
We remark that (16) defines $H_{\sigma}$ as the proximal operator of a regularisation functional $R$. However, such an $R$ exists only if $H_{\sigma}$ is a firmly nonexpansive operator — a condition not guaranteed in the learned denoising setting. In our framework, we treat this formulation as a conceptual model aligned with recent works (e.g., [10]), and do not assume the explicit existence or convexity of $R$. Our convergence results rely on the Lipschitz continuity of $H_{\sigma}$, convexity of the data fidelity term, and boundedness of the surrogate regularisation.
Differentiability of the Implicit Regulariser. The implicit regulariser $R(x)$ derived from $H_{\sigma}$ must have a Lipschitz continuous gradient. This condition can be satisfied if $H_{\sigma}(x)$ is differentiable and smooth, which follows from the construction of the KAN denoiser using B-splines. In single-shot learning, we ensure the smoothness of $H_{\sigma}$, avoiding overfitting artifacts or discontinuities introduced during adaptation to the single image. We emphasise that while (16) conceptually associates the denoiser $H_{\sigma}$ with a regularisation functional $R$, we do not assume that $R$ is differentiable. However, since the KAN denoiser $H_{\sigma}$ is constructed from smooth B-spline basis functions, it is differentiable. This differentiability enables practical use of updates in PnP-ADMM, and contributes to stable convergence behavior. Nonetheless, the differentiability of $H_{\sigma}$ does not imply the existence or differentiability of an underlying functional $R$.
@@ 2.3 Convergence of KAN-PnP with Single-Instance Prior
$\min_{x,z}f(x)+\lambda R(H_{\sigma}(z))\quad\text{subject to}\quad x=z,$
@@
    Lipschitz continuity of the KAN denoiser $H$ ensures that the denoising step is stable and does not grow arbitrarily with input perturbations. While this does not imply non-expansiveness (i.e., 1-Lipschitz continuity), convergence can still be guaranteed under this weaker assumption, as discussed in [10].
@@
    Differentiability of the regulariser $R(x)$, if required for updates, which is satisfied under the smoothness of $H_{\sigma}$.
Theorem 2.3 (Fixed-Point Convergence of KAN-PnP).
Under the assumptions:
  1. (i)
    $f:\mathbb{R}^{n}\to\mathbb{R}$ is proper, closed, and convex,
  2. (ii)
    $H_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is Lipschitz continuous with constant $L_{H}$ (as established in Theorem 1),
  3. (iii)
    The penalty parameter $\mu>L_{H}$ is constant across iterations,
the PnP-ADMM iterates $\{(x^{k},z^{k},u^{k})\}$ defined in (17) converge to a fixed point $(x^{*},z^{*},u^{*})$ satisfying:
$x^{*}=H_{\sigma}(z^{*}-u^{*}),\quad z^{*}=\arg\min_{z}\left(f(z)+\frac{\mu}{2}\|z-(x^{*}+u^{*})\|^{2}\right),\quad x^{*}=z^{*}.$
Proof Sketch.
We analyse the convergence of the KAN-PnP ADMM iterations using the fixed-point framework introduced in recent Plug-and-Play literature [10, 12]. This approach does not assume the existence of an explicit energy functional involving the denoiser, but instead analyses the behavior of the iterative updates as an operator composition acting on the optimisation variables. The denoising step $x^{k+1}=H_{\sigma}(z^{k}-u^{k})$ applies a Lipschitz continuous operator $H_{\sigma}$ with constant $L_{H}$ (as guaranteed by Theorem 1). The $z$-update step:
$z^{k+1}=\arg\min_{z}\left(f(z)+\frac{\mu}{2}\|z-(x^{k+1}+u^{k})\|^{2}\right),$
is equivalent to the proximal operator of the convex function $f$, evaluated at $x^{k+1}+u^{k}$. Since $f$ is convex and proper, the proximal operator $\text{Prox}_{f/\mu}$ is firmly nonexpansive.
The dual update $u^{k+1}=u^{k}+x^{k+1}-z^{k+1}$ is a standard ADMM step enforcing consensus between $x$ and $z$. When combining these three steps, the overall ADMM iteration can be interpreted as a fixed-point iteration of a composed operator acting on $(x,z,u)$. Under the assumption that $\mu>L_{H}$, it follows from standard results in fixed-point theory that the composition of the Lipschitz operator $H_{\sigma}$ and the firmly nonexpansive proximal operator defines an averaged operator with a contraction factor strictly less than one [10]. By the Banach fixed-point theorem, this guarantees that the sequence $\{(x^{k},z^{k},u^{k})\}$ converges to a unique fixed point $(x^{*},z^{*},u^{*})$ satisfying:
$x^{*}=H_{\sigma}(z^{*}-u^{*}),\quad z^{*}=\text{Prox}_{f/\mu}(x^{*}+u^{*}),\quad x^{*}=z^{*}.$
These conditions characterise a consistent fixed point of the algorithm, even though the objective function is implicit. This result justifies the convergence of the KAN-PnP method under mild and verifiable assumptions, without requiring knowledge of the underlying regulariser associated with $H_{\sigma}$. ∎
Remark. The convergence result in Theorem 2 adapts the fixed-point framework established in recent Plug-and-Play literature [10, 12]. While the overall structure follows these works, our analysis is specialised to the single-instance setting, where the denoiser $H_{\sigma}$ is trained directly from a noisy observation. This introduces unique architectural properties (e.g., differentiability from B-splines) and excludes any access to external datasets or explicit regularisers. Our proof reflects this structure and confirms convergence under mild assumptions on the denoiser and fidelity term.
@@ 3 Experimental Results & Discussion
In this section, we evaluate the performance of our proposed KAN-PnP framework across challenging image reconstruction tasks, including super-resolution and joint optimisation for demosaicing and deconvolution. Our approach leverages a single-instance prior, wherein the denoiser is trained directly on the observed image itself, rather than on a dataset of similar images. The goal of these experiments is to assess the effectiveness of our approach compared to other single-instance denoisers and the conventional methods, including pretrained denoisers and total variation techniques within the Plug-and-Play framework. This training scheme is designed to exploit the unique characteristics of the target image, enhancing reconstruction quality without requiring extensive external datasets. This rationale motivates our comparison with traditional methods trained in a conventional manner.
For the experimental comparison, we followed the dual resizing strategies to preprocess the image. The images are sourced with Creative Commons Licenses which are resized to $512\times 384$, and the selected data in [33, 34] were tested. The experiment were conducted on NVIDIA A100 GPU with 80GB RAM. The Plug-and-Play phase ensued with the $\bigtriangledown$-Prox toolbox [35], adopting the default settings for all tasks. Why Natural Images? While our method is applicable to diverse data types, we focus on natural images in this work for several reasons. First, natural image benchmarks (e.g., Set5, Set14) are widely used in foundational works on single-instance priors (e.g., DIP [25] and Cheng et al. [27]), enabling direct comparison with prior art. Second, these datasets offer well-understood structure and standardized metrics, which help isolate the methodological contribution of the learned prior. Third, since our aim is to investigate core behavior of single-shot Plug-and-Play priors — a relatively unexplored domain — natural images provide a valuable, interpretable setting for proof-of-concept evaluation.
@@
Evaluation Protocol. To ensure a robust comparison, we evaluated the Noise2Self pre-training approach [36] on three networks, each undergoing $100$ iterations with a learning rate of $0.01$. The FFDNet [37] model is designed with $8$ hidden layers and $64$ feature maps per layer, while the UNet [23] architecture featured a $4$-level downsampling scheme and began with $32$ feature maps in its initial convolutional layer. For consistency, the optimisation setup mirrored the configuration used in our proposed approach. We compare these results against both the Noise2Self approach and implicit neural MLP, which we denoted as INR, as well as our proposed implicit KAN method. The evaluation of the methods were measured by Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM), where PSNR and SSIM scores signal are higher when reconstruction quality improved. On Parameter Counts. While the KAN and INR architectures differ in parameter count, we do not enforce equality since they encode fundamentally different inductive biases. KANs exploit local interpolation priors, while INRs rely on global MLP representations. For this reason, we evaluate models based on reconstruction quality rather than parameter parity.
@@ 3.1 KAN-PnP in High-Stakes Image Reconstruction
Our proposed KAN-PnP framework has been extensively evaluated in two challenging inverse problem settings: super-resolution and joint optimisation for demosaicing and deconvolution. Super-resolution tests the model’s ability to enhance image detail across varying upscaling factors (2$\times$, 4$\times$ and 8$\times$), while the joint optimisation task emphasises its robustness and adaptability by tackling the intertwined challenges of demosaicing and deconvolution. The latter setting is particularly demanding due to the need to simultaneously address two operations—removing the effects of mosaic patterns while restoring the high-frequency details lost during image acquisition—requiring a balance of accuracy and computational.
@@ 4 Conclusion
In this work, we introduced KAN-PnP, an optimisation framework that integrated Kolmogorov-Arnold Networks (KANs) as denoisers within the Plug-and-Play (PnP) paradigm. KAN-PnP was specifically designed to tackle inverse problems with single-instance priors, where only a single noisy observation was available, thus eliminating the need for large datasets typically required by traditional denoising methods. We demonstrated that KANs, based on the Kolmogorov-Arnold representation theorem, effectively served as priors in these settings, offering a robust solution for denoising. We provided theoretical guarantees for KAN-PnP, proving its convergence under key assumptions. Extensive experimental results showed that KAN-PnP outperformed existing denoising methods, providing superior performance in single-shot learning scenarios with minimal data. The method exhibited reliable convergence, achieving high accuracy with fewer iterations. Comparisons with state-of-the-art approaches further highlighted KAN-PnP’s advantages in both convergence speed and solution quality, demonstrating its potential for solving real-world inverse problems with limited data. An interesting direction for future work lies in explicitly enforcing strict non-expansiveness of the KAN denoiser (i.e., ensuring $L_{H}\leq 1$) through architectural constraints. This could lead to tighter theoretical convergence guarantees and further improve the empirical stability of Plug-and-Play methods in the single-instance setting.